You wish to measure the iron content of the well water on the new property you are about to buy. You prepare a reference standard Fe3+ solution with a concentration of 7.57×10−4 M. You treat 10.0 mL of this reference with HNO3 and excess KSCN to form a red complex and dilute the reference to 50.0 mL. The diluted reference is placed in a cell with a 1.00‑cm light path. You then take 30.0 mL of the well water, treat with HNO3 and excess KSCN, and dilute to 100.0 mL. This diluted sample is placed in a variable pathlength cell. The absorbance of the reference and the sample solutions match when the pathlength is 3.36 cm. What is the concentration of iron in the well water? For each solution, the zero is set with a blank.

[Fe3+]well water=

M
Answer #1

[Fe3+]well water = 1.50 x 10-4 M

Explanation

Fe3+ reference solution concentration = 7.57 x 10-4 M

Concentration of complex = (7.57 x 10-4 M) * (10.0 mL / 50.0 mL)

Concentration of complex = 1.514 x 10-4 M

From Beer's law,

absorbance = (molar absorptivity) * (path length) * (concentration)

or A = (e) * (l) * (C)

For same absorbance value,

(l1) * (C1) = (l2) * (C2)

(1.00 cm) * (1.514 x 10-4 M) = (3.36 cm) * (C2)

C2 = (1.514 x 10-4 M) * (1.00 cm / 3.36 cm)

C2 = 4.506 x 10-5 M

Fe3+ concentration in well water = (4.506 x 10-5 M) * (100.0 mL / 30.0 mL)

Fe3+ concentration in well water = 1.50 x 10-4 M